OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (cwrw238):

first divide both sides by -3 to give |2x + 6| = 4 now, do you know what the vertical lines on either side of 2x + 6 mean?

OpenStudy (anonymous):

i have no idea :/

OpenStudy (anonymous):

Is that what shows where to split each equation?

OpenStudy (cwrw238):

it means the absolute value of 2x + 6 in other words whether 2x + 6 is negative or positive you take the positive value so 2x + 6 can be positive or negative so we write 2 equations 2x + 6 = 4 and 2x + 6 = -4 solve these 2 to get your answer

OpenStudy (cwrw238):

ok can you do that? do you understand absolute value?